# citations.bib

### v1

@@ 1 Introduction
Explaining the prediction of a machine learning model is as important as building the model itself, since it helps determine whether the model can be trusted to give meaningful predictions when deployed in real world [ribeiro2016should]. Such explanations of black-box decisions are all the more important in sensitive applications, such as medicine, finance, and law [burton2016final].
In the quest of explaining models, recent line of research has focused on developing local explanation methods with the objective to identify the degree of influence of each feature that a specific data point has on the model prediction. These include Explanation vectors [baehrens2010explain], LIME [ribeiro2016should], and Shapley values [vstrumbelj2014explaining]. When local methods are expressed as additive feature attribution methods, i.e., the feature influence linearly adds up to provide the model prediction, [lundberg2017unified] provided game theoretic results guaranteeing that Shapley values provide a unique solution to additive feature attribution. For these reasons, it has emerged as a front-runner model agnostic explanation tool.
Shapley values have found relevance in other machine learning applications too. They have been used in measuring the global sensitivity analysis where for instance they have been used to partition the coefficient of determination quantity in linear regression [song2016shapley]. More concretely, Shapley values offer a general approach of answering the following question: given a model $f$ trained on data-points with $d$ features, and evaluated on a test sample $\bm{q}\in\mathbb{R}^{d}$, how does each feature of $\bm{q}$ locally influence the final model decision $f(\bm{q})$?
Consider the value function $v:2^{[d]}\to\mathbb{R}$, where $v(S)$ depends on the output of the model on a test sample $\bm{q}$ using only the subset of features corresponding to the elements of the subset $S$ of $[d]=\{1,\ldots,d\}$. For instance, given a baseline $\bm{q}^{\textup{base}}$, we may define $v(S)=f(\bm{q}^{(S)})$ where $\bm{q}_{j}^{(S)}=\bm{q}_{j}$ if $j\in S$ and $\bm{q}_{j}^{\textup{base}}$ otherwise. [^1] There are other established ways to do this including replacing a fixed baseline with an expectation over suitable inputs or even training the model with only the features in $S$ present [chen2018shapley, lundberg2017unified]. The precise choice is not important for us, as the methods discussed in this paper work for any value function. The Shapley value $\phi_{j}^{*}$ corresponding to the $j$-th feature contribution is defined as
@@ 1.1 Fast Approximate Estimators
In general, computing Eq. 1.1 requires evaluating $v(S)$ on each of the exponentially many subsets of $[d]$. Each evaluation of $v(S)$ is costly, with the exact cost depending on the way $v(S)$ is defined. While this cost can be reduced for certain types of simple models [lundberg2018consistent], an appealing aspect of Shapley values is the potential for model-agnostic explanations.
To make Shapley values computationally tractable for arbitrary models, multiple randomized estimators have been proposed. Such methods aim to approximate the Shapley values, while using a sub-exponential number of value-function evaluations; see [castro2009polynomial, williamson2020efficient, okhrati2021multilinear, mitchell2022sampling, zhang2024fast]. Perhaps the most popular is a method called KernelSHAP, which is implemented in the widely used SHAP library [lundberg2017unified]. KernelSHAP and related estimators are the focus of this paper.
Approximate Shapley Value Estimation. In what follows, the all ones vector and zero vector are $\bm{1}$ and $\bm{0}$ respectively, and the $j$-th standard basis vector is $\bm{e}_{j}$. Given a vector $\bm{a}$, $\lVert\bm{a}\rVert$ denotes its Euclidean norm, while for a matrix $\bm{A}$, $\lVert\bm{A}\rVert$ denotes its spectral norm. The key observation [Charnes1988] used by KernelSHAP and related estimators is that the Shapley values are the solution to a certain constrained least squares problem
@@
  • $\bm{W}$ is a $(2^{d}-2)\times(2^{d}-2)$ diagonal matrix: $\bm{W}{S,S}=k(S)=(d-1)/({|S|\choose d}|S|(d-|S|))$
  • $\bm{v}$ is a $22^{d}-$ 2 length vector: $\bm{v}_{S}=v(S)-v(\emptyset)$.
As with the definition of the Shapley values Eq. 1.1, the regression formulation Eq. 1.2 requires knowledge of knowing $v(S)$ for each $2^{d}$ subsets of $[d]$. To get around this cost, KernelSHAP (randomly) subsamples and reweights the rows of Eq. 1.2, and then outputs the solution to the (much smaller) constrained regression problem. The sampling of the $S$-th row $\bm{Z}_{S}$ is done proportional to the kernel weight $k(S)$, a choice made based on the observation that the objective function $\left\lVert\bm{Z}^{\prime}\bm{\phi}-\bm{b}\right\rVert^{2}$ can be written as an expectation $\mathbb{E}[(\bm{Z}_{S}\bm{\phi}-\bm{v}_{S})^{2}]$ with respect to this sampling distribution, as explained in Section B.1. Other practical improvements such as paired-sampling and sampling without replacement are also included in the implementation of KernelSHAP in the SHAP library.
A large number of subsequent works have built on KernelSHAP [covert2020improving, lundberg2017unified, aas2021explaining, zhang2024fast, musco2025provably, jethani2021fastshap, kwon2022weightedshap, fumagalli2024kernelshap, kariyappa2024shap]. Of particular relevance to the present work are unbiased KernelSHAP [covert2020improving] and LeverageSHAP [musco2025provably] which, to the best of our knowledge, are the only extensions of KernelSHAP with theoretical convergence guarantees. The method of [covert2020improving] is an unbiased variant of KernelSHAP for which an asymptotic variance analysis is given. It was however observed that this method tends to underperform compared to the original KernelSHAP in practice. The method of [musco2025provably] is a regression-based estimator and satisfies strong non-asymptotic theoretical guarantees. Numerical experiments suggest that it may outperform KernelSHAP is essentially all settings.
High-Dimensional Estimators. Additionally, several works have specifically focused on the challenges of computing Shapley values for high-dimensional data [aas2021explaining, chen2018shapley, jethani2021fastshap, frye2020shapley, hu2024manifold, zhang2024fast]. These use parametric approaches to the computation of Shapley values; however, they require overhead model pretraining. Building on [frye2020shapley], [hu2024manifold] develop a method for high-dimensional SHAP estimation using latent features. [chen2018shapley] propose a specific approach for data structured on graphs; such approaches avoid computing SHAP for large dimensions leveraging inductive biases. Recently, [zhang2024fast] propose SimSHAP, an unbiased alternative to [jethani2021fastshap]. Methods for large language models, such as [kokalj2021bert] have recently been develop; however, no algorithm at present is tailored for high dimensional settings while providing provable guarantees on sample efficiency.
@@ 1.2 Our Contribution
In this work, we present a novel and unified framework to analyze Shapley value estimators. Using tools from randomized linear algebra, we prove non-asymptotic sample complexity guarantees on the efficient behavior of the estimators, including KernelSHAP [lundberg2017unified] and LeverageSHAP [musco2025provably]. Specifically, we identify three main contributions of the present work:
@@
These advancements promote trust in the estimation of Shapley values, enabling their usage in safety-critical applications. In Section 2, we develop the unified framework: defining the estimators and distributions in Section 2.1 and Section 2.2 respectively, and providing our main result on sample complexity guarantees in Section 2.3. In Section 3, we extensive experimental evaluation of the described estimators, comparing their performance in Section 3.1, and showcasing their effectiveness in higher dimensional settings in Section 3.2.
@@ 2 A Unified Framework for Provable Shapley Value Estimation
Towards this end, it is useful to reformulate Eq. 1.2, using a change of variable approach, in terms of an ordinary linear regression or a matrix-vector multiplication problem.
@@
Let $\bm{Q}$ be any fixed $d\times(d-1)$ matrix whose columns for an orthonormal basis for the space of vectors orthogonal to the all-ones vector (i.e. $\bm{Q}^{\mathsf{T}}\bm{Q}=\bm{I}$, $\bm{Q}^{\mathsf{T}}\bm{1}=\bm{0}$). Given $\lambda\in\mathbb{R}$, define
@@
A similar formulation of the Shapley values in terms of unconstrained regression appears in [musco2025provably]. Theorem 2.1, which is proved in Section A.3, goes beyond that of [musco2025provably] in two key ways. First, we observe that by solving the unconstrained problem explicitly, we obtain the solution as the product of a matrix $\bm{Q}\bm{U}^{\mathsf{T}}$ and vector $\bm{b}_{\lambda}$. Second, we make the observation that there is complete freedom in the choice of $\lambda\in\mathbb{R}$. Together, these advancements allow us to develop a unifying framework for providing provable guarantees for a broad class of randomized estimators which encompasses many existing estimators [covert2020improving, lundberg2017unified, aas2021explaining, zhang2024fast, musco2025provably, etc.].
@@ 2.1 Randomized Estimators Within our Framework
We frame our exposition in the context of randomized sketching, a powerful technique which has been studied for decades in randomized numerical linear algebra [woodruff2014sketching, Martinsson2020].
In the context of Shapley value estimation, a sketching matrix is a $m\times(2^{d}-2)$ matrix $\bm{S}$ where each row has exactly one nonzero entry and $\mathbb{E}[\bm{S}^{\mathsf{T}}\bm{S}]=\bm{I}$. We leave the exact choice of the distribution of $\bm{S}$ general, but discuss several natural choices in Section 2.2. Regardless of the distribution, since each of the $m$ rows of $\bm{S}$ has exactly one nonzero entry, computing $\bm{S}\bm{b}$ requires at most $m$ evaluations of $v(S)$. Thus, estimators which make use of $\bm{S}\bm{b}$ can be substantially more efficient to compute when $m\ll 2^{d}$.
Using the sketch $\bm{S}\bm{b}_{\lambda}$ (which can easily be computed from $\bm{S}\bm{b}$) in the formulations in Theorem 2.1 yields estimators based on sketched regression or on approximate matrix-vector multiplication.
@@
    Given the sketching matrix $\bm{S}$, this estimator can be computed in $O(md^{2}+mT_{v})$ time, where $T_{v}$ is the time to evaluate an entry of $\bm{b}$.
@@
We provide proofs that the estimators from [lundberg2017unified, covert2020improving, musco2025provably] fit into our framework in Appendix B. Past works, especially [covert2020improving], have used a Lagrangian framework to obtained closed-form solutions to their randomized estimators. While this is mathematically equivalent to our change of variable approach, as described in Appendix C, the expressions, which involve ratios of correlated random variables, are seemingly harder to analyze directly in the Lagrangian framework leading to previous difficulties in providing proofs of KernelSHAP [covert2020improving].
@@ 2.2 Sampling Schemes for Sketching Matrix
The choice of $\bm{S}$ plays a critical role in both the regression and matrix-vector multiplication estimators–which $m$ entries of $\bm{b}$ are observed impacts what we learn about the Shapley values. However, model-agnostic estimators cannot make strong assumptions about the structure of $\bm{b}$. The relative importance of the $i$-entry of $\bm{b}$ can be encoded in a probability distribution $\mathcal{P}$ over subsets $S\subset 2^{[d]}\setminus\{[d],\emptyset\}$. This distribution is subsequently used to generate $\bm{S}$ and hence sample the entries of $\bm{b}$ [^4] The approaches we consider only take into account the relative importance of individual rows. Other approaches (e.g. based on Determinantal Point Processes/volume sampling) take into account the relative importance of entire sets of rows. This results in stronger theoretical guarantees for general regression problems, but such distributions are harder to sample from [Dereziski2021]. In the context of Shapley value estimation, it is common to use further optimizations such as paired sampling and sampling without replacement, which we explore empirically in Section 3.
@@
    Leverage score / $\ell_{2}$-squared sampling: The LeverageSHAP method chooses sampling probabilities proportional to the statistical leverage scores of $\bm{U}$. Since $\bm{U}$ has orthonormal columns, the leverage score of the $S$th row of $\bm{U}$ coincides with the squared row-norm $\left\lVert\bm{u}_{S}\right\rVert^{2}$, which is widely used in the quantum-inspired algorithms framework [tang2019quantum]. Leverage score sampling for sketched regression satisfies strong theoretical guarantees, which [musco2025provably] use to prove guarantees about the LeverageSHAP estimator.
@@
which is the weighted geometric mean of the kernel weights and the leverage scores (see Eq. A.64 for the full expression). $\tau=1$ gives kernel weight distribution, $\tau=0$ gives leverage score sampling, while $\tau=1/2$ gives modified $\ell_{2}$ sampling.
Figure 1: Comparison of the sampling probabilities described in Section 2.2. Kernel Weights (dashed), Leverage scores (dash-dot), and our proposed modified $\ell_{2}$-weights (solid), which are the geometric mean of the Kernel Weights and Leverage scores.

### v2

# A Unified Framework for Provably Efficient Algorithms to Estimate Shapley Values

@@ 1 Introduction
Explaining the prediction of a machine learning model is as important as building the model itself, since it helps determine whether the model can be trusted to give meaningful predictions when deployed in real world [RSG16]. Such explanations of black-box decisions are all the more important in sensitive applications, such as medicine, finance, and law [Bur+16].
In the quest of explaining models, recent line of research has focused on developing local explanation methods with the objective to identify the degree of influence of each feature that a specific data point has on the model prediction. These include Explanation vectors [Bae+10], LIME [RSG16], and Shapley values [ŠK14]. When local methods are expressed as additive feature attribution methods, i.e., the feature influence linearly adds up to provide the model prediction, [LL17] provided game theoretic results guaranteeing that Shapley values provide a unique solution to additive feature attribution. For these reasons, it has emerged as a front-runner model agnostic explanation tool.
Shapley values have found relevance in other machine learning applications too. They have been used in measuring the global sensitivity analysis where for instance they have been used to partition the coefficient of determination quantity in linear regression [SNS16]. More concretely, Shapley values offer a general approach of answering the following question: given a model $f$ trained on data-points with $d$ features, and evaluated on a test sample $\bm{q}\in\mathbb{R}^{d}$, how does each feature of $\bm{q}$ locally influence the final model decision $f(\bm{q})$?
Consider the value function $v:2^{[d]}\to\mathbb{R}$, where $v(S)$ depends on the output of the model on a test sample $\bm{q}$ using only the subset of features corresponding to the elements of the subset $S$ of $d]=\{1,\ldots,d\}$. For instance, given a baseline $\bm{q}^{\textup{base}}$, we may define $v(S)=f(\bm{q}^{(S)})$ where $\bm{q}_{j}^{(S)}=\bm{q}_{j}$ if $j\in S$ and $\bm{q}_{j}^{\textup{base}}$ otherwise. [^1] There are other established ways to do this including replacing a fixed baseline with an expectation over suitable inputs or even training the model with only the features in $S$ present [CSWJ18, LL17]. The precise choice is not important for us, as the methods discussed in this paper work for any value function. The Shapley value $\phi_{j}^{*}$ corresponding to the $j$-th feature contribution is defined as
@@ 1.1 Fast Approximate Estimators
In general, computing Eq.˜1.1 requires evaluating $v(S)$ on each of the exponentially many subsets of $[d]$. Each evaluation of $v(S)$ is costly, with the exact cost depending on the way $v(S)$ is defined. While this cost can be reduced for certain types of simple models [LEL18], an appealing aspect of Shapley values is the potential for model-agnostic explanations.
To make Shapley values computationally tractable for arbitrary models, multiple randomized estimators have been proposed. Such methods aim to approximate the Shapley values, while using a sub-exponential number of value-function evaluations; see [CGT09, WF20, OL21, MCFH22, Zha+24]. Perhaps the most popular is a method called KernelSHAP, which is implemented in the widely used SHAP library [LL17]. KernelSHAP and related estimators are the focus of this paper.
Approximate Shapley Value Estimation. In what follows, the all ones vector and zero vector are $\bm{1}$ and $\bm{0}$ respectively, and the $j$-th standard basis vector is $\bm{e}_{j}$. Given a vector $\bm{a}$, $\lVert\bm{a}\rVert$ denotes its Euclidean norm, while for a matrix $\bm{A}$, $\lVert\bm{A}\rVert$ denotes its spectral norm. The key observation [CGKR88] used by KernelSHAP and related estimators is that the Shapley values are the solution to a certain constrained least squares problem
@@
  • $\bm{W}$ is a $(2^{d}-2)\times(2^{d}-2)$ diagonal matrix: $\bm{W}{S,S}=k(S)=(d-1)/({d\choose{|S|}}|S|(d-|S|))$
  • $\bm{v}$ is a $2^{d}-2$ length vector: $\bm{v}_{S}=v(S)-v(\emptyset)$.
As with the definition of the Shapley values Eq.˜1.1, the regression formulation Eq.˜1.2 requires the knowledge of $v(S)$ for each $2^{d}$ subsets of $[d]$. To get around this cost, KernelSHAP (randomly) subsamples and reweights the rows of Eq.˜1.2, and then outputs the solution to the (much smaller) constrained regression problem. The sampling of the $S$-th row $\bm{Z}_{S}$ is done proportional to the kernel weight $k(S)$, a choice made based on the observation that the objective function $\left\lVert\bm{Z}^{\prime}\bm{\phi}-\bm{b}\right\rVert^{2}$ can be written as an expectation $\mathbb{E}[(\bm{Z}_{S}\bm{\phi}-\bm{v}_{S})^{2}]$ with respect to this sampling distribution, as explained in Section˜B.1. Other practical improvements such as paired-sampling and sampling without replacement are also included in the implementation of KernelSHAP in the SHAP library.
A large number of subsequent works have built on KernelSHAP [CL20, LL17, AJL21, Zha+24, MW25, Jet+21, KZ22, Fum+24, KTLM24]. Of particular relevance to the present work are unbiased KernelSHAP [CL20] and LeverageSHAP [MW25] which, to the best of our knowledge, are the only extensions of KernelSHAP with theoretical convergence guarantees. The method of [CL20] is an unbiased variant of KernelSHAP for which an asymptotic variance analysis is given. It was however observed that this method tends to underperform compared to the original KernelSHAP in practice. The method of [MW25] is a regression-based estimator and satisfies strong non-asymptotic theoretical guarantees. Numerical experiments suggest that it may outperform KernelSHAP in most settings.
High-Dimensional Estimators. Additionally, several works have specifically focused on the challenges of computing Shapley values for high-dimensional data [AJL21, CSWJ18, Jet+21, Fry+20, HZFS24, Zha+24]. These use parametric approaches to the computation of Shapley values; however, they require overhead model pretraining. Building on [Fry+20], [HZFS24] develop a method for high-dimensional SHAP estimation using latent features. [CSWJ18] propose a specific approach for data structured on graphs; such approaches avoid computing SHAP for large dimensions leveraging inductive biases. Recently, [Zha+24] propose SimSHAP, an unbiased alternative to [Jet+21]. Methods for large language models, such as [Kok+21] have recently been develop; however, no algorithm at present is tailored for high dimensional settings while providing provable guarantees on sample efficiency.
@@ 1.2 Our Contribution
In this work, we present a novel and unified framework to analyze Shapley value estimators. Using tools from randomized linear algebra, we prove non-asymptotic sample complexity guarantees on the efficient behavior of the estimators, including KernelSHAP [LL17] and LeverageSHAP [MW25]. Specifically, we identify three main contributions of the present work:
@@
These advancements promote trust in the estimation of Shapley values, enabling their usage in safety-critical applications. In Section˜2, we develop the unified framework: defining the estimators and distributions in Section˜2.1 and Section˜2.2 respectively, and providing our main result on sample complexity guarantees in Section˜2.3. In Section˜3, we perform an extensive experimental evaluation of the described estimators, comparing their performance in Section˜3.1, and showcasing their effectiveness in higher dimensional settings in Section˜3.2.
@@ 2 A Unified Framework for Provable Shapley Value Estimation
Towards this end, it is useful to reformulate Eq.˜1.2 in terms of an ordinary linear regression or a matrix-vector multiplication problem involving a matrix with orthonormal columns. The key observation herein is that any vector $\bm{\phi}\in\mathbb{R}^{d}$ satisfying the constraint $\bm{1}^{\mathsf{T}}\bm{\phi}=v([d])-v(\emptyset)$ can be decomposed as the sum of a vector proportional to $\bm{1}$ (with proportionality constant $(v([d])-v(\emptyset))/d$) and a vector orthogonal to $\bm{1}$. By converting Eq.˜1.2 to an unconstrained problem, we will be able to more easily understand how popular Shapley value estimators can be studied through the lens of randomized numerical linear algebra.
@@
Let $\bm{Q}$ be any fixed $d\times(d-1)$ matrix whose columns form an orthonormal basis for the space of vectors orthogonal to the all-ones vector (i.e. $\bm{Q}^{\mathsf{T}}\bm{Q}=\bm{I}$, $\bm{Q}^{\mathsf{T}}\bm{1}=\bm{0}$). Given $\lambda\in\mathbb{R}$, define
@@
A similar formulation of the Shapley values in terms of unconstrained regression appears in [MW25]. Theorem˜2.1, which is proved in Section˜A.3, goes beyond that of [MW25] in two key ways. First, we observe that by solving the unconstrained problem explicitly, we obtain the solution as the product of a matrix $\bm{Q}\bm{U}^{\mathsf{T}}$ and vector $\bm{b}_{\lambda}$. Second, we make the observation that there is complete freedom in the choice of $\lambda\in\mathbb{R}$. Together, these advancements allow us to develop a unifying framework for providing provable guarantees for a broad class of randomized estimators which encompasses many existing estimators [CL20, LL17, AJL21, Zha+24, MW25, etc.].
@@ 2.1 Randomized Estimators Within our Framework
We frame our exposition in the context of randomized sketching, a powerful technique which has been studied for decades in randomized numerical linear algebra [Woo+14, MT20].
In the context of Shapley value estimation, a sketching matrix is an $m\times(2^{d}-2)$ matrix $\bm{S}$ where each row has exactly one nonzero entry and $\mathbb{E}[\bm{S}^{\mathsf{T}}\bm{S}]=\bm{I}$. We leave the exact choice of the distribution of $\bm{S}$ general, but discuss several natural choices in Section˜2.2. Regardless of the distribution, since each of the $m$ rows of $\bm{S}$ has exactly one nonzero entry, computing $\bm{S}\bm{b}$ requires at most $m$ evaluations of $v(S)$. Thus, estimators which make use of $\bm{S}\bm{b}$ can be substantially more efficient to compute when $m\ll 2^{d}$.
Using the sketch $\bm{S}\bm{b}_{\lambda}$ (which can easily be computed from $\bm{S}\bm{b}$) in the formulations in Theorem˜2.1 yields estimators based on sketched regression or on approximate matrix-vector multiplication.
@@
    Given the sketching matrix $\bm{S}$, this regression (or least squares) estimator can be computed in $O(md^{2}+mT_{v})$ time, where $T_{v}$ is the time to evaluate an entry of $\bm{b}$.
@@
We provide proofs that the estimators from [LL17, CL20, MW25] fit into our framework in Appendix˜B. Past works, especially [CL20], have used a Lagrangian framework to obtained closed-form solutions to their randomized estimators. While this is mathematically equivalent to our change of variable approach, as described in Appendix˜C, the expressions, which involve ratios of correlated random variables, are seemingly harder to analyze directly in the Lagrangian framework leading to previous difficulties in providing proofs of KernelSHAP [CL20].
@@ 2.2 Sampling Schemes for Sketching Matrix
The choice of $\bm{S}$ plays a critical role in both the regression and matrix-vector multiplication estimators–which $m$ entries of $\bm{b}$ are observed impacts what we learn about the Shapley values. However, model-agnostic estimators cannot make strong assumptions about the structure of $\bm{b}$. The relative importance of the $i$-entry of $\bm{b}$ can be encoded in a probability distribution $\mathcal{P}$ over subsets $S\subset 2^{[d]}\setminus\{[d],\emptyset\}$. This distribution is subsequently used to generate $\bm{S}$ and hence sample the entries of $\bm{b}$. [^4] The approaches we consider only take into account the relative importance of individual rows. Other approaches (e.g. based on Determinantal Point Processes/volume sampling) take into account the relative importance of entire sets of rows. This results in stronger theoretical guarantees for general regression problems, but such distributions are harder to sample from [DM21]. In the context of Shapley value estimation, it is common to use further optimizations such as paired sampling and sampling without replacement, which we explore empirically in Section˜3.
@@
    Leverage Score / $\ell_{2}$-squared Sampling: The LeverageSHAP method chooses sampling probabilities proportional to the statistical leverage scores of $\bm{U}$. Since $\bm{U}$ has orthonormal columns, the leverage score of the $S$th row of $\bm{U}$ coincides with the squared row-norm $\left\lVert\bm{u}_{S}\right\rVert^{2}$, which is widely used in the quantum-inspired algorithms framework [Tan19]. Leverage score sampling for sketched regression satisfies strong theoretical guarantees, which [MW25] use to prove guarantees about the LeverageSHAP estimator.
@@
which is the weighted geometric mean of the kernel weights and the leverage scores (see Eq.˜A.66 for the full expression). $\tau=1$ gives kernel weight distribution, $\tau=0$ gives leverage score sampling, while $\tau=1/2$ gives modified $\ell_{2}$ sampling.
Figure 1: Comparison of the sampling probabilities described in Section˜2.2. Kernel Weights (dashed), Leverage scores (dash-dot), and our proposed modified $\ell_{2}$-weights (solid), which are the geometric mean of the Kernel Weights and Leverage scores.